OpenStudy (anonymous):

w/15=3/5 a 1/9 b 1/3 c 6 d 9

OpenStudy (wendy.ivette11714):

Step 1. Cross Multiply: \[w \times(5)=(3)\times(15)\]

OpenStudy (wendy.ivette11714):

\[5 \times w = ?\] \[15 \times 3 = ?\] @jayjaysimmer

OpenStudy (anonymous):

3 times 15 is 45

OpenStudy (wendy.ivette11714):

Thats correct, now what is 5 times w?

OpenStudy (anonymous):

I Don't Know >.<

OpenStudy (wendy.ivette11714):

\[w \times 5 = 5w\]

OpenStudy (anonymous):

Oh Ok

OpenStudy (wendy.ivette11714):

Now that we multiplied those, we get: \[5w=45\]

OpenStudy (wendy.ivette11714):

Now we divide both sides by 5: \[\frac{ 5w }{ 5 }\frac{ 45 }{ 5 }\]

OpenStudy (anonymous):

45 divided by 5 is 9

OpenStudy (wendy.ivette11714):

Yes

OpenStudy (wendy.ivette11714):

So the answer would be: \[w = 9\]

OpenStudy (anonymous):

Thank You!

OpenStudy (wendy.ivette11714):

Your Welcome :)

OpenStudy (anonymous):

Can you help with another?

OpenStudy (wendy.ivette11714):

Sure, I have time for one more :)

OpenStudy (anonymous):

The ratio of red marbles to blue marbles in a jar is 3 to 8. There are 40 blue marbles in the jar. How many red marbles are in the jar? A. 5 B. 15 C. 24 D. 35

OpenStudy (anonymous):

I Think Its C 24

OpenStudy (wendy.ivette11714):

C is wrong. Here is how you solve this problem: \[3:8\] \[8 \times 5 = 40\] \[3 \times 5 = 15\] Answer: B

OpenStudy (anonymous):

Oh Ok Thank You :)

OpenStudy (anonymous):

w/15=3/5 15 is 3 times 5. w is 3 times 3 or 9.